# Random data Cauchy theory for fully nonlocal telegraph equations

### v1

@@ 1. $(\mathcal{PC})$
Current research focuses on investigating the connection between fundamental solutions of time nonlocal telegraph equations with convolution kernel $a_{1}\in(\mathcal{PC})$ and stochastic processes. Pozo and Vergara [22] established that the fundamental solution of (1.1) corresponds to the probability density function of a specific stochastic process $X(t)$. Furthermore, employing the Karamata-Feller Tauberian theory ([10, Chapter XIII]), they characterized the asymptotic behavior of the process variance across different temporal scales. Subsequently, Alegría and Pozo [2] introduced a systematic technique to generate examples where the variance $\text{Var}X(t)$ grows sublinearly or logarithmically in time, contrasting with classical diffusive scaling. Furthermore, Thang [25] established that for any given completely monotone ultraslow kernel $a_{1}$ in (1.1), there exists an induced kernel such that the mean squared displacement of the associated stochastic process for the time nonlocal telegraph equation exhibits logarithmic growth. Recently, Alegría and Pozo [3] considered a non-Markovian telegraph processes $X_{k}(t)$ whose probability density function solves the following problem
@@
numerous counterexamples demonstrate that the kernel $b\ast r_{\gamma}$ (where $r_{\gamma}$ is defined in (2.4)) fails to be completely positive. This critical limitation prevents the direct application of the subordination principle [20, Chapter 4] to derive the relevant solution operators. This issue requires urgent resolution.
@@
Here, $t>0$, $x\in\mathbb{R}^{3}$, $\gamma>0$, $\kappa>1$ and $(-\Delta)^{\frac{\beta}{2}}$ with $\beta\in(1,2]$ is the fractional Laplacian operator. The kernel $a$ is a creep function (see [20, Definnition 4.4]) such that there exists a function $b\in L_{loc}^{1}(\mathbb{R}_{+})$ satisfying $(da\ast b)(t)=1$ for $t>0$, that is, $a$ satisfies the following condition
@@ 2.2 Deterministic and probabilistic Preliminaries
We present a criterion for identifying strong $(p,q)$ type operators from [13, Theorem 1.11].
@@ 2.3 Volterra integral equations
where $s_{\gamma},r_{\gamma}\in L_{loc}^{1}(\mathbb{R}^{+})$ are the unique solution to the above, respectively, see [12, Chapter 2, Theorem 3.1]. Sometimes, $s_{\gamma}$ is also referred to as the scalar resolvent function, while $r_{\gamma}$ is called the integrated scalar resolvent function. Let $f\in L_{loc}^{p}(\mathbb{R}_{+})$ for some $p\geq 1$, then the unique solution to the following Volterra integral equation
@@
and $v_{\gamma}\in L_{loc}^{p}(\mathbb{R}_{+})$, see [12, Chapter 2, Theorem 3.5]. Furthermore, if $f\in C(\mathbb{R}_{+})$, then $v_{\gamma}\in C(\mathbb{R}_{+})$ and the previous form of $v_{\gamma}(t)$ also holds for $t=0$, see [12, Chapter 2, Theorem 3.5].
@@
Thus, $(a,b)\in(\mathcal{PC}^{\ast})$ implies that $b$ is completely positive. It should be mentioned that a completely positive function must be nonnegative, see [8, Propposition 2.1(i)]. Further equivalent conditions regarding the complete positivity of $b$ is presented in [20, Proposition 4.5].
@@
    Moreover if $b$ is nonincreasing, then
@@
The justification for (i) and (ii) can be founded in [14, Proposition 2.1]. For (iii), the proof is completely analogous to that in [21, Lemma 5.4], so we omit it. ∎
@@ Appendix A Auxiliary Lemma concerning relaxation functions
Let us first prove (i). It is shown in [20, Proposition 4.5] that $(k,l)\in(\mathcal{\mathcal{missing}}{PC^{\ast}})$ is equivalent to ${\widehat{l}(\lambda)}]^{-1}\in(\mathcal{BF})$. Further, $\lambda\mapsto e^{-{\theta}{\widehat{l}(\lambda)}^{-1}}\in(\mathcal{CM})$ for each $\theta\geq 0$, see [20, Proposition 4.2]. Applying Bernstein’s Theorem ([20, Chapter 4]), there exists a unique function $\phi(\cdot,\theta)\in BV_{loc}(\mathbb{R}_{+})$ that is nondecreasing with $\phi(0,\theta)=0$ such that
@@
According to (2.6), we obtain $\widehat{l}(\lambda)^{-1}=k_{0}\lambda+\lambda\widehat{k_{1}}(\lambda)=k_{\infty}+k_{0}\lambda+\lambda\widehat{k_{2}}(\lambda)$ for $\lambda>0$. It is stated in [20, Proposition 4.3] that $\inf\limits_{\lambda>0}\widehat{l}(\lambda)^{-1}=k_{\infty}$. Thus, when $\gamma\in\{z\in\mathbb{C}:Re(z)\geq-k_{\infty}\}$, Integrating (A.2) yields
@@
The lemma below serves to address the $L_{t}L_{x}$ estimates to problem (1.4). The proof is a direct adaptation of that for [23, Lemma 2, Lemma 3]. The key modification lies in that we relax the condition $(\mathcal{PC})$ to the complete positive condition.
@@ Appendix B New completely positive kernels
Since $d_{\delta_{1}},d_{\delta_{2},\gamma}\in L_{loc}^{1}(\mathbb{R}_{+})$, [12, Theorem 2.2(i)] shows that $h_{\delta_{1},\delta_{2},\gamma}:=(d_{\delta_{1}}\ast d_{\delta_{2},\gamma})(t)$ is well-defined for $t>0$ and $h_{\delta_{1},\delta_{2},\gamma}\in L_{loc}^{1}(\mathbb{R}_{+})$. Let us to prove that $h_{\delta_{1},\delta_{2},\gamma}$ is completely positive. Applying the Laplace transform to $h_{\delta_{1},\delta_{2},\gamma}$ yields
@@
Due to the assumption of $\delta_{1},\delta_{2}$ and [24, Proposition 7.13] we obtain that $\lambda\mapsto\lambda^{1-\delta_{1}}(\lambda+\gamma)^{1-\delta_{2}}$ is a Bernstein function. Meanwhile, since the kernel $l$ is completely positive, [20, Proposition 4.5] yields that $\widehat{l}(\lambda)]^{-1}$ is a Bernstein function. Note that $\mathcal{BF}\circ\mathcal{BF}\in\mathcal{BF}$([20, Proposition 4.2]). Then (B.2) shows that $\widehat{h}_{\delta_{1},\delta_{2},\gamma}(\lambda)^{-1}\in(\mathcal{B}F)$, which implies $h_{\delta_{1},\delta_{2},\gamma}$ is completely positive by [20, proposition 4.5].
@@ Appendix C Estimates for the fractional Laplacian
Employing the approach developed in [4, Theorem], we can readily generalize to derive the following key result.
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm it|\xi|^{\frac{\beta}{2}}}\tilde{u}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim t^{\frac{2s}{\beta}-\frac{6}{\beta}(\frac{1}{p}-\frac{1}{p^{\prime}})}\|u\|_{H^{s,p}}\text{ ~for~ }t>0.$ (C.2)
@@
Define $Q(\xi):=|\xi|^{\frac{\beta}{2}}$ with $\xi\in\mathbb{R}^{3}$. A straightforward computation yields the Hessian matrix
@@
$\text{rank}(H_{Q})=\begin{cases}2\text{ ~for~ }\beta=2,\ \xi\neq 0,\\ 3\text{ ~for~ }\beta\in(1,2),\ \xi\neq 0.\end{cases}$
Denote $\varphi=\text{rank}(H_{Q})$ and $Q_{1}(\xi)=|\xi|^{s}$ with $\xi\in\mathbb{R}^{3}$. We now proceed to prove
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\tilde{u}(\xi)\right)\right\|_{L^{\infty}}\lesssim 2^{(3-s-\frac{\varphi}{4}\beta)j}\|u\|_{L^{1}},\quad j\in\mathbb{N}_{+}.$ (C.3)
@@
$\begin{aligned} \left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i2^{\frac{\beta j}{2}}Q(\xi)}\varrho(\xi)\right)\right\|_{L^{\infty}}\leq L(\varrho/Q_{1},Q)2^{-\frac{\beta\varphi j}{4}}\text{ ~for~ }j\in\mathbb{N}_{+},\end{aligned}$ (C.4)
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\right)\right\|_{L^{\infty}}\lesssim 2^{(3-s-\frac{\varphi}{4}\beta)j},\quad j\in\mathbb{N}_{+}.$
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\tilde{u}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim 2^{((3-\frac{\beta}{4}\varphi)(\frac{1}{p}-\frac{1}{p^{\prime}})-s)j}\|u\|_{L^{p}}\leq\|u\|_{L^{p}},\quad j\in\mathbb{N}_{+}.$ (C.6)
Here, we use the fact $(3-\frac{\beta}{4}\varphi)(\frac{1}{p}-\frac{1}{p^{\prime}})-s\leq 0$, provided the assumption (C.1). On the other hand, note that $1<p\leq\frac{6}{3+s}$ implies $\frac{1}{p}-\frac{1}{p^{\prime}}\in[\frac{s}{3},1)$. Let $\psi$ satisfy $\frac{1}{\psi}=\frac{1}{p}-\frac{1}{p^{\prime}}$, then $\psi\in(1,\frac{3}{s}]$. Since $\varrho_{0}\in C_{c}^{\infty}(\mathbb{R}^{3})$, it follows directly that
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-\frac{\beta}{2}}\sin(|\xi|^{\frac{\beta}{2}})\varrho_{j}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim 2^{((3-\frac{\beta}{4}\varphi)(\frac{1}{p}-\frac{1}{p^{\prime}})-\frac{\beta}{2})j}\|u\|_{L^{p}}\leq\|u\|_{L^{p}},\quad j\in\mathbb{N}_{+}$
by (C.6). Here, we use the fact that $(3-\frac{\beta}{4}\varphi)(\frac{1}{p}-\frac{1}{p^{\prime}})-\frac{\beta}{2}\leq 0$, provided the assumption (C.10). The following proof follows the same argument as in the proof subsequent to (3.7) in Lemma C.1, and is therefore omitted here. ∎
@@
For $\beta=2$, Proposition C.2 reduces to [4, Theorem] with $n=3$.

### v3

@@ 1. $(\mathcal{PC})$
Current research focuses on investigating the connection between fundamental solutions of time nonlocal telegraph equations with convolution kernel $a_{1}\in(\mathcal{PC})$ and stochastic processes. Pozo and Vergara [27] established that the fundamental solution of (1.1) corresponds to the probability density function of a specific stochastic process $X(t)$. Furthermore, employing the Karamata-Feller Tauberian theory ([12, Chapter XIII]), they characterized the asymptotic behavior of the process variance across different temporal scales. Subsequently, Alegría and Pozo [2] introduced a systematic technique to generate examples where the variance $\text{Var}X(t)$ grows sublinearly or logarithmically in time, contrasting with classical diffusive scaling. Furthermore, Thang [31] established that for any given completely monotone ultraslow kernel $a_{1}$ in (1.1), there exists an induced kernel such that the mean squared displacement of the associated stochastic process for the time nonlocal telegraph equation exhibits logarithmic growth. Recently, Alegría and Pozo [3] considered a non-Markovian telegraph processes $X_{k}(t)$ whose probability density function solves the following problem
@@
numerous counterexamples demonstrate that the kernel $b\ast r_{\gamma}$ (where $r_{\gamma}$ is defined in (2.4)) fails to be completely positive. This critical limitation prevents the direct application of the subordination principle [25, Chapter 4] to derive the relevant solution operators. This issue requires urgent resolution.
@@
Here, $t>0$, $x\in\mathbb{R}^{3}$, $\gamma>0$, $\kappa>1$ and $(-\Delta)^{\frac{\beta}{2}}$ with $\beta\in(1,2]$ is the fractional Laplacian operator. The kernel $a$ is a creep function (see [25, Definnition 4.4]) such that there exists a function $b\in L_{loc}^{1}(\mathbb{R}_{+})$ satisfying $(da\ast b)(t)=1$ for $t>0$, that is, $a$ satisfies the following condition
@@ 2.2 Deterministic and probabilistic Preliminaries
We present a criterion for identifying strong $(p,q)$ type operators from [15, Theorem 1.11].
@@ 2.3 Volterra integral equations
where $s_{\gamma},r_{\gamma}\in L_{loc}^{1}(\mathbb{R}^{+})$ are the unique solution to the above, respectively, see [14, Chapter 2, Theorem 3.1]. Sometimes, $s_{\gamma}$ is also referred to as the scalar resolvent function, while $r_{\gamma}$ is called the integrated scalar resolvent function. Let $f\in L_{loc}^{p}(\mathbb{R}_{+})$ for some $p\geq 1$, then the unique solution to the following Volterra integral equation
@@
and $v_{\gamma}\in L_{loc}^{p}(\mathbb{R}_{+})$, see [14, Chapter 2, Theorem 3.5]. Furthermore, if $f\in C(\mathbb{R}_{+})$, then $v_{\gamma}\in C(\mathbb{R}_{+})$ and the previous form of $v_{\gamma}(t)$ also holds for $t=0$, see [14, Chapter 2, Theorem 3.5].
@@
Thus, $(a,b)\in(\mathcal{PC}^{\ast})$ implies that $b$ is completely positive. It should be mentioned that a completely positive function must be nonnegative, see [8, Propposition 2.1(i)]. Further equivalent conditions regarding the complete positivity of $b$ is presented in [25, Proposition 4.5].
@@
    Moreover if $l$ is nonincreasing, then
@@
The justification for (i) and (ii) can be founded in [16, Proposition 2.1]. For (iii), the proof is completely analogous to that in [26, Lemma 5.4], so we omit it. ∎
@@ Appendix A Auxiliary Lemma concerning relaxation functions
Let us first prove (i). It is shown in [25, Proposition 4.5] that $(k,l)\in(\mathcal{\mathcal{}}{PC^{\ast}})$ is equivalent to ${\widehat{l}(\lambda)}]^{-1}\in(\mathcal{BF})$. Further, $\lambda\mapsto e^{-{\theta}{\widehat{l}(\lambda)}^{-1}}\in(\mathcal{CM})$ for each $\theta\geq 0$, see [25, Proposition 4.2]. Applying Bernstein’s Theorem ([25, Chapter 4]), there exists a unique function $\phi(\cdot,\theta)\in BV_{loc}(\mathbb{R}_{+})$ that is nondecreasing with $\phi(0,\theta)=0$ such that
@@
According to (2.6), we obtain $\widehat{l}(\lambda)^{-1}=k_{0}\lambda+\lambda\widehat{k_{1}}(\lambda)=k_{\infty}+k_{0}\lambda+\lambda\widehat{k_{2}}(\lambda)$ for $\lambda>0$. It is stated in [25, Proposition 4.3] that $\inf\limits_{\lambda>0}\widehat{l}(\lambda)^{-1}=k_{\infty}$. Thus, when $\gamma\in\{z\in\mathbb{C}:Re(z)\geq-k_{\infty}\}$, Integrating (A.2) yields
@@
The lemma below serves to address the $L_{t}L_{x}$ estimates to problem (1.4). The proof is a direct adaptation of that for [28, Lemma 2, Lemma 3]. The key modification lies in that we relax the condition $(\mathcal{PC})$ to the complete positive condition.
@@ Appendix B New completely positive kernels
Since $d_{\delta_{1}},d_{\delta_{2},\gamma}\in L_{loc}^{1}(\mathbb{R}_{+})$, [14, Theorem 2.2(i)] shows that $h_{\delta_{1},\delta_{2},\gamma}:=(d_{\delta_{1}}\ast d_{\delta_{2},\gamma})(t)$ is well-defined for $t>0$ and $h_{\delta_{1},\delta_{2},\gamma}\in L_{loc}^{1}(\mathbb{R}_{+})$. Let us to prove that $h_{\delta_{1},\delta_{2},\gamma}$ is completely positive. Applying the Laplace transform to $h_{\delta_{1},\delta_{2},\gamma}$ yields
@@
Due to the assumption of $\delta_{1},\delta_{2}$ and [30, Proposition 7.13] we obtain that $\lambda\mapsto\lambda^{1-\delta_{1}}(\lambda+\gamma)^{1-\delta_{2}}$ is a Bernstein function. Meanwhile, since the kernel $l$ is completely positive, [25, Proposition 4.5] yields that $\widehat{l}(\lambda)]^{-1}$ is a Bernstein function. Note that $\mathcal{BF}\circ\mathcal{BF}\in\mathcal{BF}$([25, Proposition 4.2]). Then (B.2) shows that $\widehat{h}_{\delta_{1},\delta_{2},\gamma}(\lambda)^{-1}\in(\mathcal{B}F)$, which implies $h_{\delta_{1},\delta_{2},\gamma}$ is completely positive by [25, proposition 4.5].
@@ Appendix C Estimates for the fractional Laplacian
Early research on the decay estimates for the operator $\mathcal{F}^{-1}(|\xi|^{-s}e^{it|\xi|^{\alpha}}\mathcal{F})$ with $s>0$ can be traced back to [23] by Pecher. He considered the case where $\alpha$ is an integer. In fact, employing the approach of Pecher, we can readily generalize to derive the case where $\alpha\in(\frac{1}{2},1)$. For other related research, see [24, 29, 21, 11, 10].
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm it|\xi|^{\frac{\beta}{2}}}\tilde{u}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim t^{\frac{2s}{\beta}-\frac{6}{\beta}(\frac{1}{p}-\frac{1}{p^{\prime}})}\|u\|_{L^{p}}\text{ ~for~ }t>0.$ (C.2)
@@
The case $\beta=2$ has been established in [23, Theorem 2.2](cf. also [24, Theorem 1.1]). It therefore suffices to consider the scenario where $\beta\in(1,2)$. Define $Q(\xi):=|\xi|^{\frac{\beta}{2}}$ with $\xi\in\mathbb{R}^{3}$. A straightforward computation yields the Hessian matrix
@@
$\text{rank}(H_{Q})=3\text{ ~for~ }\beta\in(1,2),\ \xi\neq 0.$
Denote $Q_{1}(\xi)=|\xi|^{s}$ with $\xi\in\mathbb{R}^{3}$. We now proceed to prove
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\tilde{u}(\xi)\right)\right\|_{L^{\infty}}\lesssim 2^{(3-s-\frac{3}{4}\beta)j}\|u\|_{L^{1}},\quad j\in\mathbb{N}_{+}.$ (C.3)
@@
$\begin{aligned} \left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i2^{\frac{\beta j}{2}}Q(\xi)}\varrho(\xi)\right)\right\|_{L^{\infty}}\leq L(\varrho/Q_{1},Q)2^{-\frac{3\beta j}{4}}\text{ ~for~ }j\in\mathbb{N}_{+},\end{aligned}$ (C.4)
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\right)\right\|_{L^{\infty}}\lesssim 2^{(3-s-\frac{3}{4}\beta)j},\quad j\in\mathbb{N}_{+}.$
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-s}e^{\pm i|\xi|^{\frac{\beta}{2}}}\varrho_{j}(\xi)\tilde{u}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim 2^{((3-\frac{3\beta}{4})(\frac{1}{p}-\frac{1}{p^{\prime}})-s)j}\|u\|_{L^{p}}\leq\|u\|_{L^{p}},\quad j\in\mathbb{N}_{+}.$ (C.6)
Here, we use the fact $(3-\frac{3\beta}{4})(\frac{1}{p}-\frac{1}{p^{\prime}})-s\leq 0$, provided the assumption (C.1). On the other hand, note that $1<p\leq\frac{6}{3+s}$ implies $\frac{1}{p}-\frac{1}{p^{\prime}}\in[\frac{s}{3},1)$. Let $\psi$ satisfy $\frac{1}{\psi}=\frac{1}{p}-\frac{1}{p^{\prime}}$, then $\psi\in(1,\frac{3}{s}]$. Since $\varrho_{0}\in C_{c}^{\infty}(\mathbb{R}^{3})$, it follows directly that
@@
$\left\|\mathcal{F}^{-1}\left(|\xi|^{-\frac{\beta}{2}}\sin(|\xi|^{\frac{\beta}{2}})\varrho_{j}(\xi)\right)\right\|_{L^{p^{\prime}}}\lesssim 2^{((3-\frac{3\beta}{4})(\frac{1}{p}-\frac{1}{p^{\prime}})-\frac{\beta}{2})j}\|u\|_{L^{p}}\leq\|u\|_{L^{p}},\quad j\in\mathbb{N}_{+}$
by (C.6). Here, we use the fact that $(3-\frac{3\beta}{4})(\frac{1}{p}-\frac{1}{p^{\prime}})-\frac{\beta}{2}\leq 0$, provided the assumption (C.10). The following proof follows the same argument as in the proof subsequent to (3.7) in Lemma C.1, and is therefore omitted here. ∎
@@
For $\beta=2$, Proposition C.2 reduces to a special case of [4, Theorem].